Brett and Julie headed north at 8 a.m. By noon Brett was 80 miles ahead of Julie. What was Brett's speed if Julie's speed was 30 miles per hour?
Found 2 solutions by ikleyn, greenestamps:
Answer by ikleyn(52810)

.

Let x be the Brett' speed.


Then you have this equation 


    4x = 4*30 + 80


saying that in 4 hours traveled time Brett covered the distance of 4x miles, while Julie covered 4*30 miles 

and Brett's distance was 80 miles longer then that of Julie.


From the equation,


    4x = 120 + 80 = 200,

     x            = 200/4 = 50 miles per hour.    ANSWER

Solved.


Answer by greenestamps(13203)



Brett covered 80 miles more than Julie in 4 hours.

So his rate was 80/4 = 20 miles per hour faster than Julie's.

Since Julie's rate was 30mph, Brett's rate was 30+20 = 50mph.